### v2

# IntGrad MT: Eliciting LLMs’ Machine Translation Capabilities with Sentence Interpolation and Gradual MT

Abstract
Recent Large Language Models (LLMs) have demonstrated strong performance in translation without needing to be finetuned on additional parallel corpora. However, they still underperform for low-resource language pairs. Previous works have focused on mitigating this issue by leveraging relevant few-shot examples or external resources such as dictionaries or grammar books, making models heavily reliant on these nonparametric sources of information. In this paper, we propose a novel method named IntGrad MT that focuses on fully exploiting an LLM’s inherent translation capability. IntGrad MT achieves this by constructing a chain of few-shot examples, each consisting of a source sentence and the model’s own translation, that rise incrementally in difficulty. IntGrad MT employs two techniques: Sentence Interpolation, which generates a sequence of sentences that gradually change from an easy sentence to translate to a difficult one, and Gradual MT, which sequentially translates this chain using translations of earlier sentences as few-shot examples for the translation of subsequent ones. With this approach, we observe a substantial enhancement in the xCOMET scores of various LLMs for multiple languages, especially in low-resource languages such as Hindi(8.26), Swahili(7.10), Bengali(6.97) and Marathi(13.03). Our approach presents a practical way of enhancing LLMs’ performance without extra training.
1 Introduction
Recent Large Language Models (LLMs) have shown strong performance in translation tasks without the need for fine-tuning on specific parallel datasets. Previous studies have demonstrated that LLMs’ translation capabilities are reliable in most use cases, particularly when the source and target language are high-resource languages (Zhu et al., 2024; Robinson et al., 2023; Jiao et al., 2023). However, because LLMs require training on large corpora, they still face challenges when translating low-resource languages that are not sufficiently represented in the training corpora.(Stap & Araabi, 2023; Robinson et al., 2023; Enis & Hopkins, 2024).
Previous research has attempted to address these challenges by leveraging the in-context-learning capabilities of large language models (LLMs), particularly through the use of external knowledge such as few-shot examples or dictionaries during inference. However, relevant examples are not always guaranteed to be available, and constructing such external knowledge sources can be costly. A potential solution is to reduce reliance on external sources altogether.
In this paper, we examine whether we can improve LLMs’ translation capabilities without relying on external knowledge. We aim to answer this question by considering a simple fact: the machine translation task can be defined as a mapping between two (sub)spaces for the source and target language. In the source language space, there are regions where the model performs well in translation and regions where it does not. The key idea is that if we can gradually expand the areas where the model performs well by feeding it neighboring examples, we can enhance its translation capabilities in the areas where it performs poorly.
We propose IntGrad MT to achieve this by connecting sentences from the regions where the LLM performs well in translation and regions where it does not. IntGrad MT consists of two key techniques: Sentence Interpolation and Gradual MT. Sentence Interpolation is a prompting technique that generates a sequence of sentences gradually transitioning from one to another. Gradual MT is a technique in which the model iteratively translates a list of sentences, using its previous translations as few-shot examples for subsequent sentences. Through sentence interpolation, we first establish a pathway to the sentences in regions where the model performs poorly, then gradually expand the area where the model can excel, utilizing its in-context learning capabilities. The key concepts are illustrated in Figure 1.
Figure 1: Figurative illustration of IntGrad MT. Machine translation task can be defined as a mapping between two (sub)spaces for the source and target language. In the source language space, there are regions where the model performs well in translation and regions where it does not. IntGrad MT expand the areas where the model performs well by feeding LLM with neighboring examples, eventually reaching the areas where it performs poorly.
We test the effectiveness of IntGrad MT by applying it to four different LLMs — GPT-3.5 Turbo, Mistral Nemo Instruct, Llama 3.1 70B Instruct, and Llama 3.1 8B Instruct — across seven target languages: German (De), Chinese (Zh), Hindi (Hi), Korean (Ko), Swahili (Sw), Marathi (Mr), and Bengali (Bn). Our results show consistently large performance gains, particularly in low-resource languages.
Additionally, we conduct an ablation study on English-to-Korean translation using ChatGPT to determine the optimal settings in relation to the following three questions:
  • How should we select the start sentences for interpolation?
  • How should we aggregate multiple translation results from different interpolation paths?
  • To what sentences IntGradMT can be effective?
Our contributions are summarized as follows:
  • We propose IntGrad MT, a novel method for machine translation that leverages two key methods: Sentence Interpolation and Gradual MT.
  • We test the effectiveness of IntGrad MT with various LLMs and across different target languages.
  • We conduct an ablation study to identify the optimal configuration.
  • We analyze the geometrical characteristics of sentence interpolation.
2 Related Work
2.1 Enhancing LLMs’ Translation Capabilities without Fine-tuning
Modern LLMs show high capabilities in translation tasks in high- resource languages, but not in low-resource languages (Jiao et al., 2023; Stap & Araabi, 2023; Zhu et al., 2024; Enis & Hopkins, 2024). There are several works that have focused on enhancing LLMs’ translation capabilities without additional fine-tuning. A primary strategy involves leveraging LLMs’ ability to learn from demonstrations or descriptions (Brown et al., 2020; Wei et al., 2022). Studies have explored selecting appropriate exemplars for few-shot learning and demonstrating linguistic knowledge (Agrawal et al., 2022; Vilar et al., 2023; Zhang et al., 2024), or augmenting LLMs with chains of multilingual dictionaries (Lu et al., 2024). Besides providing a demonstration or description, choosing the right temperature or prompting strategy has also been examined (Peng et al., 2023). Similar to previous research, our method aims to improve LLMs’ MT capabilities without fine-tuning, focusing instead on eliciting the models’ inherent capabilities with sentence interpolation and gradual MT rather than providing them with few-shot examples or external knowledge.
2.2 Self-Demonstration
Manually generating appropriate exemplars for in-context learning can be resource-intensive. To address this challenge, previous studies have explored enabling models to generate their own few-shot examples for tasks such as classification(Lyu et al., 2023; Kim et al., 2022) or other reasoning tasks(Zhang et al., 2023; Li et al., 2024). Our work is aligned with these efforts, as it also focuses on generating the model’s own few-shot examples. However, none of these approaches have tried to create examples such as ours, since our approach aims to produce the tailored example by gradually expanding the example set with similar yet distinct examples.
3 Methods
IntGrad MT consists of two main steps: Sentence Interpolation (§3.1) and Gradual MT (§3.2). We first introduce these two steps and present how we combine them (§3.3).
3.1 Sentence Interpolation
We propose sentence interpolation, which is a prompting technique asks model to create a list of sentences that gradually change from start sentence to end sentence.
For example, following is the prompt that we use in our experiments:
{shadedquotation}
I will give you two sentences. Can you gradually change the first sentence to make it exactly the same as the second sentence? Just give me the sentences and don’t provide additional comments.
Sentence1: $\langle Sentence1\rangle$
Sentence2: $\langle Sentence2\rangle$
The objective of this technique is to generate a list of sentences where each sentence is distinct, yet not excessively different from its adjacent sentences. By prompting the LLM to autonomously create the list, rather than mechanically altering sentences, we can obtain natural sentences that are suitable for use as in-context learning examples. We call this list of interpolated sentences the interpolation path.
In practice, we utilized three few-shot examples from GPT-4(OpenAI et al., 2024) to control the output format. See Appendices A and B for these examples and sample interpolation paths.
3.2 Gradual MT
Gradual MT is a prompting technique that lets an LLM utilize its previous translations as a prompt. This approach sequentially processes a list of sentences, translating each one while using the previous translation results as few-shot examples for the current sentence. An illustration of the Gradual MT process is shown in Figure 2. Gradual MT can effectively guide LLMs in accurately translating unfamiliar sentences when combined with sentence interpolation.
Figure 2: Illustration of Gradual MT. Gradual MT iteratively processes a list of sentences, translating each one while using the previous translation results as few-shot examples for the current sentence.
3.3 Overall Method
IntGrad MT combines sentence interpolation and gradual MT. The illustration of the algorithm can be seen in Figure 3.
Figure 3: Illustration of the IntGrad MT algorithm. IntGrad MT combines Sentence Interpolation with Gradual MT. In step 1, $k$ start sentences are selected from a pre-created start sentence pool. In step 2, start sentences are interpolated to the end sentence, creating $k$ individual interpolation paths. Each path is then processed through Gradual MT, generating translation results for each sentence. In step 3, MT results from each interpolation path are aggregated to produce a single output MT. In step 4, the sentences to which IntGrad MT is applied are selected.
Step 0: Start Sentence Pool Creation.
Before applying IntGrad MT, the start sentence pool must be created before. This pool consists of sentences that the LLM can already translate accurately. It is crucial to ensure that the source sentences included in this pool have been tested and verified, confirming the LLM can generate high-quality translations in a zero-shot setting. We refer to the source sentences in the start sentence pool as start sentences.
Step 1: Start Sentence Selection
IntGrad MT begins with selecting $k$ start sentences from the start sentence pool by calculating the similarity with the source sentence that the LLM is trying to translate, which we call an end sentence. Measuring similarity between sentences can be done in various ways. In this paper, we utilized SBERT similarity(Reimers & Gurevych, 2019) as a primary metric. Details about start sentence selection strategy can be found in Section 6.1.1.
Step 2: Sentence Interpolation & Gradual MT.
After selecting $k$ start sentences, sentence interpolation is performed between each start sentence and the corresponding end sentence, creating $k$ individual interpolation paths. The paths are then processed through the Gradual MT, generating translation results for each sentence.
Step 3: MT Results Aggregation.
After generating a list of translation, we proceed to aggregate the final translation from each path. If a single start sentence was selected in Step 1, this step is skipped. There are several methods for aggregating translations; in our approach, we input all the translation results into the LLM once again as few-shot examples to generate the final translation. A detailed explanation of this aggregation strategy can be found in Section 6.1.2. After aggregation, we obtain the final translation, which we refer to as the output.
Step 4: Output Selection.
The final step is to select the sentences to which we apply IntGrad MT. There are two motivations for not applying the method to all sentences. First, if we decide which sentences will undergo all the steps of IntGrad MT before executing them, we can substantially reduce the computing cost. Second, even if we don’t select sentences beforehand, it is still meaningful to selectively use the outcome only when it is better than the zero-shot translation result by comparing reference-free QE model scores. We tested three different strategies in our ablation study (§6.1.3): ‘Thr.’, ‘Delta’, and ‘Thr.+Delta’. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only if the QE score is below a certain threshold. ‘Delta’ chooses outputs only if they outperform zero-shot translations. ‘Thr. +Delta’ combines these two strategies.
4 Experiment
4.1 Setup
Models.
For translation we use four different LLMs: ChatGPT(GPT-3.5-Turbo-0125) [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, Mistral Nemo(Mistral-Nemo-Instruct-2407) [^2] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407/ and two different sized Llama 3.1 Instruct models(Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct)(Dubey et al., 2024). ChatGPT is accessed via OpenAI’s API, and the others are run locally. For sentence interpolation, we employ Qwen2-72b-Instruct (Yang et al., 2024) with quantization. See Appendices C and D for settings for translation and interpolation. For output selection, we utilize a reference-free QE model CometKiwi (Rei et al., 2023) to avoid peeking at the gold translations. CometKiwi produces a Direct Assessment (DA) score, which rates translation quality on a scale from 0 to 100, normalized to a range of 0 to 1. Lastly, we used all-mpnet-base-v2333https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for SBERT sismilarity calculation.
Target Languages.
We fixed English as the source language. The target languages tested in the experiments are German (De), Chinese (Zh), Korean (Ko), Hindi (Hi), Swahili (Sw), Bengali (Bn), and Marathi (Mr). Based on Joshi et al. (2020)’s 6 scale taxonomy of language resource level, we classify German and Chinese as high resource languages, Korean and Hindi as mid resource, and the rest as low resource. For the Llama 3.1 models, we conduct tests only on German and Hindi, as those models do not support other languages. Mistral Nemo Instruct does not officially support Ko, Hi, Sw, Bn, and Mr, but we conduct experiments on those languages nevertheless, as it has some capability to generate them.
Dataset.
We use the FLORES-200 benchmark dataset(Team et al., 2022) for validation and evaluation. We first utilize the dev split of the dataset to create start sentence pool. During evaluation, to test the effect of the output selection strategy, we selected 10% of the test portion of the dataset to set the DA score threshold. The remaining 90% is used to assess the overall performance on the dataset.
Start Sentence Pool Creation.
The initial sentence pool is created by translating source sentences from the dev split of the FLORES-200 dataset using a zero-shot approach. Each source sentence is translated five times and evaluated with xCOMET (Guerreiro et al., 2024). xCOMET predicts DA score normalized to a range of 0 to 1. The most frequently occurring translation is selected as the representative translation. If no translation is repeated, the one with a score closest to the average is chosen. After selecting each representative translation, the top 100 translation pairs with the highest DA scores are selected.
Baselines.
We compare the outcomes with zero-shot MT and three-shot MT results. To assess the pure impact of our method, we use the same start sentences for interpolation as in the few-shot examples.
Evaluation.
We use the DA score evaluated by xCOMET as the primary metric for our evaluation, multiplying it by 100 for readability. Additionally, we employ MetricX (Juraska et al., 2023) to obtain an MQM score, which assesses translation errors on a scale from 0 to 25, with lower scores indicating better quality. Lastly, we also check the DA score evaluated by CometKiwi.
5 Results
We first conducted various ablation settings for English-Korean translation using ChatGPT to identify the optimal configuration such as start selection(step 1) and result aggregation(step 4). Subsequently, this configuration was applied to other IntGrad MT experiments utilizing different LLMs and languages. A detailed explanation of the ablation study is provided in the next section. In this section, we report xCOMET results of IntGrad MT on various LLMs and target languages in Table 1. Similar results based on MetricX and CometKiwi are presented in Appendix E.
As shown in Table 1, IntGrad MT outperforms the baselines in all models, whether large or small, and languages of diverse resource levels, with no degeneration that often occurs when reusing the model’s own output. This indicates that LLMs possess hidden multilingual and linguistic knowledge that can be exploited for machine translation with the correct techniques. Compared to the zero-shot baseline, IntGrad MT increases the xCOMET score by a maximum of 3.03, 8.26, and 13.03 points in high, mid, and low resource languages, respectively. It incurs the largest improvements for low- and mid- resource languages, suggesting that it holds the potential to improve the quality of translation for these languages that LLMs are currently unable to translate well. IntGrad MT also outperforms the 3 shot baselines by wide margins in low resource languages. Mistral Nemo’s xCOMET scores improve by 5.15, 4.61, 4.42 for Swahili, Bengali, Marathi, respectively. This supports the advantage of our method in formulating tailored prompts using the example translation pairs over simply listing them.
Table 1: xCOMET scores of IntGrad MT with the best ablation setting for different LLMs and target languages. The strategy used includes selecting the starting sentences by sorting with SBERT similarity and tree edit distance, aggregating MT results by feeding all gradual MT results into the LLM, and selecting the output MT only if its CometKiwi score is higher than the zero-shot translation. The results of setting the output selection strategy as Thr.+Delta, which selects the output MT only if the CometKiwi score of the zero-shot MT is below a certain threshold and the output MT’s score is higher than the zero-shot score, are presented below the main results as supplementary information. Zero-shot and three-shot are the baselines, with the three-shot examples selected using the same starting sentence selection method as IntGrad MT. Scores are multiplied by 100 for readability, and the first two highest scores for each MT model and language are highlighted in bold and underlined. Llama translation results for languages other than De and Hi are not available, as Llama does not support those languages. The numbers following the up arrows indicate the delta with respect to the zero-shot baseline.
High Resource Mid Resource Low Resource
MT Model Method DE ZH HI KO SW BN MR
GPT 3.5 0 shot 97.63 91.30 71.89 89.48 81.23 68.73 44.53
3 shot 97.83↑0.20 91.65↑0.35 72.86↑0.97 90.26↑0.78 81.43↑0.20 69.28↑0.55 45.97↑1.43
IntGradDelta 98.04↑0.40 92.42↑1.12 77.54↑5.65 92.54↑3.06 84.03↑2.80 75.60↑6.86 53.57↑9.04
IntGradThr.+Delta 97.95↑0.31 91.54↑3.03 77.27↑5.38 92.08↑2.60 83.73↑2.50 74.78↑6.05 52.31↑7.78
Llama 3.1 70B 0 shot 97.33 - 79.41 - - - -
3 shot 97.48↑0.16 - 81.19↑1.78 - - - -
IntGradDelta 97.98↑0.66 - 84.45↑5.04 - - - -
IntGradThr.+Delta 97.33↑0.00 - 84.35↑4.94 - - - -
Mistral Nemo 0 shot 96.70 88.01 66.78 81.43 38.97 71.89 43.99
3 shot 97.57↑0.86 90.81↑2.79 69.71↑2.93 88.59↑7.16 40.93↑1.95 74.25↑2.36 52.60↑8.61
IntGradDelta 97.88↑1.18 91.04↑3.03 74.19↑7.41 89.38↑7.94 46.08↑7.10 78.86↑6.97 57.02↑13.03
IntGradThr.+Delta 97.84↑1.13 91.04↑3.03 73.47↑6.69 89.31↑7.88 44.53↑5.56 78.67↑6.78 56.79↑12.81
Llama 3.1 8B 0 shot 94.99 - 69.93 - - - -
3 shot 96.27↑1.27 - 73.07↑3.14 - - - -
IntGradDelta 96.42↑1.42 - 78.18↑8.26 - - - -
IntGradThr.+Delta 95.37↑0.38 - 77.33↑7.41 - - - -
For a more balanced examination of how much gains IntGrad MT brings relative to the original performance, we compared the error reduction rate across three different metrics, xCOMET, CometKiwi, and MetricX, in Figure 4. Since the DA score reflects how ‘good’ the translation is, the error for the DA score was calculated by subtracting the score from 100 ($100-DA$). The MQM score from MetricX was used directly, as it inherently represents the error. IntGrad MT consistently reduces errors in all metrics for all languages with only one exception. Compared to Table 1 the improvements to the high resource languages are more clearly visible.
Figure 4: Error reduction rate for each target language and model, calculated as the ratio of reduced error relative to the error of zero-shot translation results. Error scores from xCOMET and CometKiwi were obtained by subtracting their values from 100, while MetricX’s score was used directly.
6 Analysis
6.1 Ablation Study
To find the optimal combination of strategies for the IntGrad MT, we conducted an ablation study on En-Ko translation task using ChatGPT in terms of four different dimensions: Start sentence selection strategy, Number of start sentences, MT result aggregation strategy, and Output selection strategy. In this section, we analyze the effects of each strategy by averaging the QE scores. Full results with all different combination of ablation settings can be found in Appendix G.
6.1.1 Start Sentence Selection Strategy
Based on the intuition that start sentences that are similar to the source sentence will be helpful, we utilized three different metrics — SBERT similarity(Reimers & Gurevych, 2019), Levenshtein distance(Levenshtein & others, 1966), and Tree edit distance(Zhang & Shasha, 1989) — to calculate similarity between sentences, and combined them in theee different ways. The first way is Sort, which sorts sentences by multiple metrics with varying priorities. The second way, Filter, initially selects the top 10 sentences based on SBERT similarity, then sorts the selection using the other metrics. The third way, Tops, picks the top sentence based on the highest similarity scores from each metric. As shown in Table 2, the selection strategies that produced the highest scores differed across aggregation methods. We chose sorting by SBERT similarity and then tree edit distance (Sort(S-T)), as it yielded the highest average scores.
Table 2: Averaged scores with different start selection strategies, start sentence numbers and aggregation strategies applied to EN-KO translation task. ‘Sort’, ‘Filter’, and ‘Tops’ denote the start sentence selection methods. Each letter in parentheses represents a similarity metric, with the order indicating the priority of these metrics. ‘L’ stands for Levenshtein distance, ‘T’ stands for tree edit distance, and ‘S’ stands for SBERT similarity. The highest average values for each axis are highlighted in bold. The highest values for each start sentence number and aggregation strategy are highlighted in underline.
Start Selection Strategy Start Sents.Num & Aggregation Strategy Average
1 (N/A) 3 Poll 3 Prompt
Filter(T-L) 91.29 91.22 91.72 91.41
Filter(L-T) 91.67 91.23 91.59 91.50
Sort(L-S) 91.35 91.22 91.57 91.38
Sort(T-S) 90.95 91.09 91.65 91.23
Sort(L-T-S) 91.19 91.21 91.50 91.30
Sort(T-L-S) 91.19 91.00 91.93 91.37
Sort(S-T) 91.50 91.39 91.83 91.57
Tops - 91.23 91.69 91.46
Average 91.31 91.20 91.69
6.1.2 Number of Start Sentences & MT Aggregation Strategy
When selecting the start sentences, we need to decide whether to use more than one start sentence. If we choose more than one, we must aggregate each translation result generated by Gradual MT. We investigated two distinct strategies for aggregating the results of Gradual MT. The first method, referred to as Polling, selects the MT result with the highest number of duplicates, drawing inspiration from prior research on self-consistency (Wang et al., 2023). If no duplicates are found, a result is selected randomly. The second method, Prompting, involves feeding all Gradual MT results into the LLM as few-shot examples to generate the final MT output. As shown in Table 2, the prompting strategy outperforms polling by 0.49 points. Polling is even worse than using a single start sentence.
6.1.3 Output Selection Strategy.
We tested three strategies for output selection. The first strategy, ‘Thr.’, aims to minimize costs by evaluating zero-shot translation results with a QE model and applying IntGrad MT only when the QE score falls below a certain threshold. The second strategy, ‘Delta’, prioritizes maximizing performance by applying IntGrad MT first and using its output only if the QE score exceeds that of the zero-shot translation. The third strategy, ‘Thr. + Delta’, combines the first two: applying IntGrad MT when the zero-shot translation’s QE score is below a threshold, and only if IntGrad MT’s score is higher. We employed CometKiwi, a reference-free QE model, to implement these strategies. To compare them, we analyzed xCOMET scores and score changes of selected outputs, applying the optimal strategies for start selection (‘Sort(S-T)’) and aggregation (‘Prompting’). As shown in Table 3, all strategies improved overall performance, with ‘Delta’ achieving a notable gain of over 1 point in QE scores compared to zero-shot MT. The results also indicate that ‘Thr. + Delta’ reduces interpolation by more than half while maintaining comparable performance to ‘Delta’, offering an effective compromise between computational efficiency and translation quality. Results with every combination of ablation strategies are shown in Table 7 in Appendix G.
Table 3: Average xCOMET scores and score changes of selected outputs when applying the optimal strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. ‘All’ selects every output. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only when the CometKiwi score is below a threshold. ‘Delta’ selects outputs only if they outperform zero-shot translations. ‘Thr. + Delta’ combines these two strategies. We also report the number of end sentences for which interpolation and Gradual MT is executed (‘No. of Interpolated End Sents’) and the number of end sentences for which the IntGrad MT output is selected over zero-shot translation (‘No. of Selected Outputs’). The results show that ‘Thr. + Delta’ reduces the number of interpolation by more than half while maintaining nearly the same translation performance.
Output Adoption Strategy Avg. Score Avg. Score Change of Selected Outputs No. of Interpolated End Sents(%) No. of Selected Outputs (%)
Zeroshot 89.48 - - -
All 91.22 1.74 911 (100%) 911 (100%)
Thr. 91.50 4.49 410 (45%) 410 (45%)
Delta 92.54 5.61 911 (100%) 497 (54%)
Thr. + Delta 92.08 8.50 410 (45%) 279 (31%)
Based on the results of the ablation study, we concluded that the optimal combination of strategies is as follows: for Start Sentence Selection, sorting with SBERT similarity followed by tree edit distance proved most effective(Sort(S-T); for the Number of Start Sentences, using 3 was optimal; for Aggregation, the best method was prompting; and for Output Selection, both ‘Delta’ and ‘Thr.+Delta’ were optimal. The chosen setting for the start sentence selection strategy differs from the configuration that yields the highest xCOMET score among all ablations (92.69 in Table 6), which sorts sentences by tree edit distance, followed by Levenshtein distance, and then SBERT similarity(Sort(T-L-S)). However, after considering the averaged effects across multiple tests, we opted to prioritize the method that selects start sentences by initially sorting with SBERT similarity.
6.2 Sentence Interpolation Analysis
We conducted an analysis of the interpolated sentences used in the ablation study to verify whether they effectively interpolate between the start and end sentences. The total number of sentences in interpolated paths is 12,729.
Sentence Interpolation Error Rate.
We defined that a situation where the first and last sentences in the interpolation path are not exactly the same as the start and end sentences as an error, and executed zero-shot translation for that end sentence. Sentence interpolation was successfully executed without error in 93.68% of cases.
Progresses of Interpolation.
We examined whether LLMs genuinely interpolate between start and end sentences or simply generate random sentences. To assess this, we use SBERT to embed the interpolated sentences and calculate their Euclidean distances from the end sentence. If these distances generally decrease, it indicates successful interpolation. To mesure this, we defined progress as follows:
$\text{progress}=d_{n-1,e}-d_{n,e}$
where $d_{n,e}$ denotes the euclidian distance between nth interpolated sentence in each interpolation path and end sentence.
As shown in Figure 5, the progress of interpolated sentences is generally positive, indicating that the interpolation effectively bridges the two sentences. The average progress for all interpolated sentences is 0.14, with a standard deviation of 0.23.
We also conducted a qualitative analysis to identify patterns in interpolation. First, we sampled 100 interpolation paths for En-Ko translation containing more than three sentences and obtained sentence embeddings using SBERT. Next, we plotted each interpolation path on a 2D plane using Principal Component Analysis (PCA). Despite the variety of patterns, we identified four notable ones, which we named based on their shapes: Arc, Triangle, Zig-Zag, and Leap. Figure 5 illustrates these four major patterns. The Arc pattern was the most common. In this pattern, the sentences gradually shift toward the end sentence, following an arc-shaped trajectory. While the Arc pattern presents a relatively smooth shape, the Zig-Zag and Triangle patterns exhibit more spiky trajectories, though the sentences still progress toward the end sentence. The final pattern, Leap, shows a single large shift toward the end sentence, without the gradual progression seen in the other patterns. Although the Leap pattern might initially be perceived as a “bad interpolation,” closer examination reveals that similar intervals also appear in other patterns. Overall, our qualitative analysis showed that, regardless of the pattern, sentence interpolation typically shifts the start sentence toward the end sentence. Additional plot samples can be found in Appendix F.
Figure 5: Distribution of progress in sentence interpolation (left) and 2D scatter plots showing four major patterns of interpolation paths (right). We analyzed interpolation paths from the En-Ko ablation study. Progress indicates how much each interpolated sentence moves closer to the target sentence, with an average progress of 0.14. Scatter plots are projected from SBERT embeddings onto a 2D plane using PCA. The X and Y axes of each plot represent the first and second principal components, respectively. Arrows in each plot show the trajectory of sentence shifts from the start sentence (blue-colored dot) to the end sentence (yellow-colored dot).
7 Conclusion
In this paper, we proposed IntGrad MT, a novel method to enhance the machine translation capabilities of various LLMs. IntGrad MT leverages sentence interpolation to guide models, eliciting stronger translation performance. Experimental results across various models and languages demonstrate that our approach consistently improves translation quality, particularly in low-resource languages, achieving meaningful gains in performance metrics. Our approach is practical in that it does not require extra training and does not conflict with previous methods that utilize other kinds of prompting techniques.
8 Limitations
Despite its success, IntGrad MT introduces some computational overhead, particularly in scenarios involving multiple start sentences or large-scale models. Future work could focus on optimizing computational efficiency or extending the approach to specialized domains and additional low-resource language pairs. Moreover, since sentence interpolation did not perform well in languages other than English, even with GPT-4 (OpenAI et al., 2024), we had to limit our focus to cases where English was the source language. Exploring better prompting techniques to interpolate non-English sentences is a potential future research direction.
Appendix A Few-shot Examples for Sentence Interpolation
In this study, we gave three examples to the interpolation model. We arbitrarily picked start and end sentences from the New York Times news articles.
A.1 Example 1
{shadedquotation}
Sentence 1: Surprisingly often, animals show up uninvited at sporting events. Sometimes, it gets a little weird.
Sentence 2: D. Cohen tried to parry attacks on his honesty as Todd Blanche, Donald J. Trump’s lawyer, tried to destroy his credibility with jurors.
Interpolation path
  1. 1.
    Surprisingly often, animals show up uninvited at sporting events. Sometimes, it gets a little weird.
  2. 2.
    Often, unexpected events occur during live events, and sometimes these can get quite weird.
  3. 3.
    At public events, such as sports or trials, unexpected guests often cause disturbances, leading to weird situations.
  4. 4.
    In public spectacles, like sports or courtrooms, unexpected participants can disrupt the normal proceedings in strange ways.
  5. 5.
    During public hearings, like in court, surprising accusations and defenses can emerge, often causing odd disruptions.
  6. 6.
    In courtroom battles, unexpected arguments and witnesses can often turn the proceedings weird as each side tries to undermine the other.
  7. 7.
    In legal battles, lawyers frequently confront each other with surprising tactics to challenge credibility, which can make the proceedings seem strange.
  8. 8.
    Michael D. Cohen, during his trial, encountered surprising tactics by Todd Blanche, Donald J. Trump’s lawyer, aiming to disrupt his credibility in unusual ways.
  9. 9.
    Michael D. Cohen tried to parry attacks on his honesty as Todd Blanche, Donald J. Trump’s lawyer, tried to destroy his credibility with jurors.
A.2 Example 2
{shadedquotation}
Sentence 1: Georgia’s Parliament overrode a presidential veto to give final approval on Tuesday evening to a contentious bill that has plunged the country into a political crisis and threatened to derail the pro-Western aspirations of many Georgians in favor of closer ties with Russia.
Sentence 2: At that time, the buildings were abandoned and inaccessible to the public, and the famous ferry that shuttled immigrants from the island to Manhattan sank in a storm in 1968.
Interpolation Path
  1. 1.
    Georgia’s Parliament overrode a presidential veto to give final approval on Tuesday evening to a contentious bill that has plunged the country into a political crisis and threatened to derail the pro-Western aspirations of many Georgians in favor of closer ties with Russia.
  2. 2.
    The decision made by Georgia’s Parliament on Tuesday evening, to override a presidential veto and approve a controversial bill, has sparked a political crisis in the country and raised concerns about the impact on the country’s pro-Western stance and relations with Russia.
  3. 3.
    The controversial bill approved by Georgia’s Parliament on Tuesday, despite a presidential veto, has caused a political crisis and stirred debates about the country’s Western alignment and its relation with Russia.
  4. 4.
    The unilateral move by Georgia’s Parliament on Tuesday to push through a divisive bill despite opposition from the president has led to a political crisis and raised concerns about the nation’s Western affiliations and ties with Russia.
  5. 5.
    The recent decision by Georgia’s Parliament to defy the presidential veto and pass a controversial bill has created a political turmoil, questioning the country’s Western orientation and relationship with Russia.
  6. 6.
    Georgia’s Parliament’s decision to override the presidential veto and approve the contentious bill has caused a political crisis in the country, prompting discussions on the impact on its Western ties and relations with Russia.
  7. 7.
    Georgia’s Parliament’s act of rejecting the presidential veto and endorsing the controversial bill on Tuesday has resulted in a political crisis, with implications for the country’s pro-Western stance and connections with Russia.
  8. 8.
    Georgia’s Parliament moved forward with finalizing a contentious bill on Tuesday, overriding the presidential veto, which has triggered a political crisis and raised concerns about the nation’s Western aspirations and relationship with Russia.
  9. 9.
    At that time, the buildings were abandoned and inaccessible to the public, and the famous ferry that shuttled immigrants from the island to Manhattan sank in a storm in 1968.
A.3 Example 3
{shadedquotation}
Sentence 1: The cake has become one of Jamaica’s most popular desserts, often found in various stores.
Sentence 2: Some plans might cover only generic versions of certain contraceptives, but patients are still entitled to coverage of a specific product that their providers deem medically necessary.
Interpolation Path
  1. 1.
    The cake has become one of Jamaica’s most popular desserts, often found in various stores.
  2. 2.
    Some plans might only cover basic versions of popular desserts like the cake, while others are more comprehensive.
  3. 3.
    In many cases, plans might cover only generic versions of popular products, but specific varieties might still be available.
  4. 4.
    Some plans might cover only generic versions of certain products, with exceptions allowed for specific needs.
  5. 5.
    Some plans might cover only generic versions of certain healthcare products, but specific options are still accessible if necessary.
  6. 6.
    Some plans might cover only generic versions of certain contraceptives, but patients can still get specific types if deemed necessary.
  7. 7.
    Some plans might cover only generic versions of certain contraceptives, but patients are still entitled to coverage of a specific product that their providers deem medically necessary.
Appendix B Sample Sentence Interpolation Path
Below is a sample of the interpolation path generated by Qwen2-72b-Intstruct during the experiment.
B.1 Sample 1
{shadedquotation}
Prompt
Sentence1: I am happy that there are people willing to support me.
Sentence2: “We now have 4-month-old mice that are non-diabetic that used to be diabetic,” he added.
Output
  1. 1.
    I am happy that there are people willing to support me.
  2. 2.
    It’s good to know that support is available when needed.
  3. 3.
    It’s reassuring to see support being provided in various situations.
  4. 4.
    He mentioned seeing positive changes in different scenarios.
  5. 5.
    “We have seen positive transformations in our subjects,” he added.
  6. 6.
    “We now have improved conditions in our test subjects,” he added.
  7. 7.
    “We now have 4-month-old mice that are non-diabetic that used to be diabetic,” he added.
B.2 Sample 2
{shadedquotation}
Prompt
Sentence1: The death toll is at least 15, a figure which is expected to rise.
Sentence2: The moisture on your hands will react with the outer layers, which will feel funny and form a sort of shell.
Output
  1. 1.
    The death toll is at least 15, a figure which is expected to rise.
  2. 2.
    The number affected is at least 15, an amount which could increase due to underlying factors.
  3. 3.
    An initial count shows at least 15 impacted, with this figure potentially growing as outer layers are peeled back.
  4. 4.
    A preliminary assessment reveals at least 15 affected individuals, with the outer layers revealing more as they are examined.
  5. 5.
    The moisture on your hands will interact with the outer layers, which may reveal more than initially thought.
  6. 6.
    The moisture on your hands will react with the outer layers, which will feel peculiar and begin to peel away.
  7. 7.
    The moisture on your hands will react with the outer layers, which will feel funny and form a sort of shell.
Appendix C Prompts and Settings for Translation
ChatGPT
ChatGPT(gpt-3.5-turbo-0125) was used via API for translation with the same prompt from OpenAI’s official documentation. [^4] https://platform.openai.com/docs/examples Temperature and top_p were set to 0.3 and 1, respectively. The actual prompt is as follows:
{shadedquotation}
System: You will be provided with a sentence in English, and your task is to translate it into $\langle$ Target Language $\rangle$.
User: $\langle$ Sentence $\rangle$
Llama-3.1 70B & 8B
[…]
User: $\langle$ Sentence $\rangle$
Mistral-Nemo-Instruct-2407
[…]
Sentence: $\langle$ Sentence $\rangle$
Appendix D Settings for Sentence Interpolation
For sentence interpolation Qwen2-72B-Instruct model was used. It was run on one A6000 GPU with 4-bit quantization using transformers library. Temperature and top_p were set to 0.6 and 0.9, respectively.
Appendix E Test Results with Different Metrics
We utilized MetricX (Juraska et al., 2023) and CometKiwi (Rei et al., 2023) to obtain supplementary metrics for evaluation. MetricX and CometKiwi are fine-tuned to predict MQM and DA scores, respectively. Tables 5 and 4 show the results from each model.
Table 4: CometKiwi results of IntGrad MT with the best ablation settings for different target languages.
High Resource Mid Resource Low Resource
MT Model Method DE ZH HI KO SW BN MR
GPT 3.5 0 shot 86.12 85.84 69.12 87.73 83.24 67.93 56.23
3 shot 86.47↑0.34 86.57↑0.72 69.53↑0.41 88.53↑0.79 83.06↓0.18 68.80↑0.86 58.19↑1.96
IntGradDelta 87.43↑1.31 87.61↑1.76 74.33↑5.21 90.92↑3.19 86.15↑2.91 75.60↑7.67 66.21↑9.99
IntGradThr.+Delta 87.15↑1.03 86.35↑0.50 74.14↑5.02 90.66↑2.93 85.98↑2.74 74.94↑7.01 65.60↑9.37
Llama 3.1 70B 0 shot 84.97 - 74.55 - - - -
3 shot 85.38↑0.41 - 75.02↑0.47 - - - -
IntGradDelta 87.15↑2.17 - 78.02↑3.47 - - - -
IntGradThr.+Delta 85.06↑0.09 - 77.93↑3.38 - - - -
Mistral Nemo 0 shot 83.32 82.90 59.72 78.41 42.72 62.03 52.53
3 shot 85.22↑1.89 85.40↑2.49 66.62↑6.90 87.20↑8.79 47.04↑4.32 72.80↑10.78 63.60↑11.07
IntGradDelta 87.09↑3.77 87.13↑4.22 70.61↑10.89 89.35↑10.94 54.46↑11.75 77.91↑15.89 67.89↑15.36
IntGradThr.+Delta 87.00↑3.67 87.13↑4.22 70.22↑10.50 89.32↑10.91 53.39↑10.67 77.76↑15.74 67.77↑15.24
Llama 3.1 8B 0 shot 80.06 - 66.71 - - - -
3 shot 82.11↑2.05 - 69.25↑2.55 - - - -
IntGradDelta 84.71↑4.65 - 73.64↑6.93 - - - -
IntGradThr.+Delta 80.63↑0.56 - 73.24↑6.53 - - - -
Table 5: MetricX results of IntGrad MT with the best ablation settings for different target languages.
High Resource Mid Resource Low Resource
MT Model Method DE ZH HI KO SW BN MR
GPT 3.5 0 shot 0.59 1.10 1.22 0.60 1.32 2.28 2.49
3 shot 0.56↓0.03 1.03↓0.07 1.18↓0.04 0.54↓0.05 1.30↓0.03 2.17↓0.11 2.23↓0.26
IntGradDelta 0.54↓0.04 1.03↓0.07 0.91↓0.31 0.44↓0.16 1.18↓0.15 1.55↓0.73 1.42↓1.07
IntGradThr.+Delta 0.55↓0.04 1.09↓0.02 0.92↓0.30 0.44↓0.16 1.18↓0.14 1.59↓0.69 1.45↓1.04
Llama 3.1 70B 0 shot 0.66 - 0.91 - - - -
3 shot 0.63↓0.02 - 0.80↓0.11 - - - -
IntGradDelta 0.58↓0.08 - 0.68↓0.23 - - - -
IntGradThr.+Delta $0$ 0.68↓0.23 - - - -
Mistral Nemo 0 shot 0.69 1.28 1.66 1.21 7.21 1.91 2.63
3 shot 0.58↓0.11 1.11↓0.16 1.25↓0.42 0.58↓0.63 6.58↓0.63 1.60↓0.31 1.60↓1.03
IntGradDelta 0.54↓0.15 1.09↓0.19 1.11↓0.56 0.54↓0.67 5.45↓1.76 1.13↓0.78 1.26↓1.37
IntGradThr.+Delta 0.55↓0.14 1.09↓0.19 1.13↓0.53 0.54↓0.66 5.56↓1.65 1.13↓0.78 1.26↓1.37
Llama 3.1 8B 0 shot 1.09 - 1.33 - - - -
3 shot 0.81↓0.28 - 1.07↓0.25 - - - -
IntGradDelta 1.13↑0.04 - 0.90↓0.43 - - - -
IntGradThr.+Delta 1.02↓0.07 - 0.92↓0.41 - - - -
Appendix F Samples of interpolation paths
We sampled 100 interpolation paths for En-Ko translation containing more than three sentences and obtained sentence embeddings using SBERT. Each interpolation path was then plotted on a 2D plane using Principal Component Analysis (PCA). Figure 6 shows 24 samples from these 100 paths.
Figure 6: 24 samples of 2D scatter plots. The plots are projected from SBERT embeddings onto a 2D plane using PCA. The X and Y axes of each plot represent the first and second principal components, respectively. Arrows in each plot show the trajectory of sentence shifts from the start sentence (blue-colored dot) to the end sentence (yellow-colored dot).
Appendix G Full Results of Ablation Study
Table 6 shows the results for every combination of strategies that we explored in the ablation study(§6.1) with En-Ko translation. Table 7 presents the average results and changes in xCOMET scores for each combination of start selection strategies, the number of start sentences, and MT aggregation strategies.
Table 6: Full results on ablation study with En-Ko translation task. All scores are measured with DA score by xCOMET. ‘Sort’, ‘Filter’, and ‘Tops’ denote the start sentence selection methods. Each letter in parentheses denotes a similarity metric, with the order indicating the priority of the metrics. ‘L’ stands for Levenshtein distance, ‘T’ stands for tree edit distance, and ‘S’ stands for SBERT similarity. The highest results for each start selection strategy are highlighted in bold, while the second-highest results are underlined.
Start Selection Filtering Aggregation Baseline (3shot)
None Poll Prompt
Sort (S-T) All 90.69 90.42 91.22 90.26
Thr. 91.31 91.14 91.50
Delta 92.29 92.18 92.54
Thr. + Delta 91.70 91.84 92.08
Sort (T-S) All 89.93 90.02 90.93 90.57
Thr. 90.31 90.71 91.24
Delta 91.81 91.97 92.47
Thr. + Delta 91.74 91.64 91.98
Sort (L-S) All 90.38 90.27 90.86 90.27
Thr. 91.10 90.95 91.12
Delta 92.08 91.96 92.41
Thr. + Delta 91.83 91.68 91.91
Sort (T-L-S) All 90.17 89.99 91.32 90.44
Thr. 90.54 90.29 91.53
Delta 92.06 91.90 92.69
Thr. + Delta 92.00 91.81 92.20
Sort (L-T-S) All 90.13 90.11 90.77 90.38
Thr. 90.69 90.85 91.13
Delta 91.99 92.05 92.24
Thr. + Delta 91.93 91.82 91.86
SBERT Filter + Sort (T-L) All 90.36 90.23 91.07 90.35
Thr. 91.04 90.88 91.30
Delta 92.07 92.10 92.53
Thr. + Delta 91.70 91.69 92.00
SBERT Filter + Sort (L-T) All 90.92 90.55 91.19 90.06
Thr. 91.46 90.98 91.23
Delta 92.37 92.29 92.61
Thr. + Delta 91.95 91.12 91.34
Tops All - 90.12 90.98 90.10
Thr. - 90.65 91.30
Delta - 92.10 92.46
Thr. + Delta - 92.06 92.01
Table 7: Average xCOMET scores and score changes of selected outputs when applying every combination of strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. ‘All’ selects every output. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only when the CometKiwi score is below a threshold. ‘Delta’ selects outputs only if they outperform zero-shot translations. ‘Thr. + Delta’ combines these two strategies. We also report the number of end sentences for which interpolation and Gradual MT is executed (‘No. of Interpolated End Sents’) and the number of end sentences for which the IntGrad MT output is selected over zero-shot translation (‘No. of Selected Outputs’). The results show that ‘Thr. + Delta’ reduces the number of interpolation by more than half while maintaining nearly the same translation performance.
Output Adoption Strategy Avg. Score Avg. Score Change of Selected Outputs No. of Interpolated Sents (%) No. of Selected Outputs (%)
Zeroshot 89.48 - - -
All 90.55 1.06 21864 (100%) 21864 (100%)
Thr. 91.01 3.70 10468 (48%) 10468 (48%)
Delta 92.21 5.30 21864 (100%) 11270 (52%)
Thr. + Delta 91.81 7.98 10468 (48%) 6664 (30%)